Given the integer function f(x) = int(x/2),

A. Find f(1.2)
B. f(-1.8)

Answer by ikleyn(53996)

.
Given the integer function f(x) = int(x/2),
(A) Find f(1.2)
(B) f(-1.8)
~~~~~~~~~~~~~~~~~~~~~

Use the definition of the function int(x).

The definition says: 


    +----------------------------------------------------------+
    |   function int(x) is defined for all real values of x.   |
    |                                                          |
    |   If x is an integer number, then f(x) is equal to x.    |
    |                                                          |
    |   If x is not an integer number, then int(x) is equal    |
    |       to the closest to x lesser integer value.          |
    +----------------------------------------------------------+


Therefore,  f(1.2)  = int(1.2/2)  = int(0.6)  =  0.     ANSWER to (A)

            f(-1.8) = int(-1.8/2) = int(-0.9) = -1.     ANSWER to (B)

Solved.


///////////////////////


Comment from student: For example, if x = 2 (an integer), then f(x) = x. So, for f(x) = int(x) where x is 2, the answer is 2.
Decimal number numbers are rational numbers. This means we must round off.
So, for f(x) = int(x) when x = 2.5, the answer is 3. Is this what you are saying?


My response: It is not what I am saying. Your treatment of my post is incorrect.
You write something, which is out of logic, and try to attribute it to me.
To get my writing in correct way, read it again, as it is written,
and do not try to attribute me what I did not write.